OpenStudy (anonymous):

The volume, V, of a cone with EQUAL radius, r, and height, h, is given by: V= pie r^2h/3 . Determine the radius and height of a cone of sand having a volume of 25000 cm^3. Round your answer to the nearest cm.

OpenStudy (johnweldon1993):

\[\large V = \pi r^2 \frac{h}{3}\] We know that r = h ...so we can just make that 'h' another 'r' ...so now we have \[\large V = \pi r^2 \frac{r}{3}\] We are given that V = 25000 so plug that in for 'V' \[\large 25000 = \pi r^2 \frac{r}{3}\] Alright...now....we have \[\large 25000 = \frac{\pi r^2 \times r}{3}\] What is r^2 times r? well r^3 of course... \[\large 25000 = \frac{\pi r^3}{3}\] Now we just solve for 'r' can you do that?

OpenStudy (anonymous):

yes r= (25000/pie) 1/3

OpenStudy (johnweldon1993):

Hmm not quite... \[\large 25000 = \frac{\pi r^3}{3}\] The first thing we want to do..is get r^3 by itself....we can do that by multiplying both sides of the equation by 3...because notice...it will eliminate the fraction on the right side \[\large 75000 = \pi r^3\] Now..we want to divide both sides by \(\large \pi \) to isolate r^3 completely \[\large r^3 = \frac{75000}{\pi}\] Make sense?

OpenStudy (anonymous):

ya now i understand sorry lol but how would you put that into your calculator

OpenStudy (johnweldon1993):

Well you would look for the \(\large \pi \) symbol And for NOW...just put 75000 / \(\large \pi\)

OpenStudy (anonymous):

so r = 23873.24

OpenStudy (johnweldon1993):

Almost... Remember I just wanted you to focus on the right side of the equation for now.... We actually had \[\large r^3 = \frac{75000}{\pi}\] We now know that the right side = 23873.24 So \[\large r^3 = 23873.24\] Notice that r^3....that means " r cubed " it's like "r squared" but instead of 2 as an exponent....it's 3

OpenStudy (johnweldon1993):

That means...to solve this...we need to take the cubed root of both sides of this equation... \[\huge \sqrt[3]{r^3} = \sqrt[3]{23873.24}\] \[\huge r = \sqrt[3]{23873.24}\] \[\large r = ?\]

OpenStudy (anonymous):

r= 28.79

OpenStudy (johnweldon1993):

Perfect! but remember to round to the nearest cm as the question states...so r = 29 !!

OpenStudy (anonymous):

thank you soo much i perfectly understand this now :)

OpenStudy (johnweldon1993):

I'm glad :))

OpenStudy (anonymous):

:)